OpenStudy (anonymous):

find the domain and range of g(x)\ g(x) = 3 SQRT(x+1) +2

OpenStudy (anonymous):

@modphysnoob :( I cant figure it out without a graph

OpenStudy (anonymous):

\[3\sqrt{x+1}+2\]?

OpenStudy (anonymous):

the only thing restricting your domain is what is under the radical.

OpenStudy (anonymous):

squrt(a) a must be greater than or equal to 0 a in our case = x+1 x+1>=0 x>=-1

OpenStudy (anonymous):

range 2 to infinity